OpenStudy (anonymous):

will medal and fan What is -12 divided by 2/3

OpenStudy (anonymous):

@igreen

OpenStudy (jhannybean):

\[\frac{-12}{\dfrac{2}{3}}\]\[\frac{-12\cdot 3}{2} =~?\]

OpenStudy (jhannybean):

No direct answers @kittenlover731

OpenStudy (jhannybean):

That is incorrect.

OpenStudy (jhannybean):

\[-12 \ne \frac{1}{12}\]

OpenStudy (jhannybean):

\[12^{-1} \implies \frac{1}{12}\]

OpenStudy (anonymous):

the link I gave you is really helpful for all fraction equations

OpenStudy (jhannybean):

All you have to remember is:\[\frac{a}{\dfrac{b}{c}} = \frac{a\cdot c}{b}\]